Asiamah was given 25% trade discount on an article listed GHC750.00. He sold it with a cash discount of 5%. Calculate Asiamah percentage profit
THANK YOU

Answer by Theo(13342)

if i understand what you are saying correctly, then:

the list price is 750.
the trade discount is 25%.
therefore you buy the item for 750 minus .25 * 750 = 750 minus 187.5 = 562.5
you then sell the item at 5% discount off the list price.
you sell it for 750 minus .05 * 750 = 750 minus 37.5 = 712.5
your cost was 562.5 and you sold it for 712.5
your profit is 712.5 minus 562.5 = 150.
your percent profit is 150 / 562.5 * 100 = 26.67% rounded to 2 decimal digits.